# A Survey on Adversarial Contention Resolution

### v1

@@ 5. Conclusion
We have surveyed the vast literature on contention resolution by delineating results based on key model assumptions, metrics, and algorithm-design choices. The problem of contention resolution remains an active topic of research and here we discuss some potential directions for future research.
Machine learning (ML) may have an important role to play in the design of contention-resolution algorithms. As discussed in Section 3.3.2, Gilbert et al. (gilbert2021contention, ) demonstrate the impact of system-size predictions, and a natural question is whether there are other useful predictions that give improved performance. For example, can predictions about whether a slot will be noisy (or jammed) yield improved robustness?
In many wireless systems, collisions have significant cost. For example, in WiFi networks, the entire packet must be sent before the sender learns of success or failure (due to a collision). Under code division multiple access modulation schemes (recall (bender:coded, )), multiple packets may be sent simultaneously; therefore, a collision ruins what could otherwise be multiple successful transmissions in a slot. In mobile settings, failure to transmit a packet may mean incurring a significant delay before the sender is next within range of the intended receiver. Therefore, modeling collision cost is another angle that presents some interesting open problems in the area of contention resolution.
@@
Energy efficiency is likely to remain an important consideration in wireless systems. The limits of trading off between robustness to worst-case noise, energy efficiency, and latency remain unknown. For example, what is the optimal energy efficiency when the channel does not admit ternary feedback and/or participants cannot communicate information to each other via $\Theta(\log n)$-sized messages? Can we achieve asymptotically-optimal latency against adversarial jamming using only, say, ${\texttt{poly}({\log n})}$ number of channel-access attempts per packet?

### v3

@@ 5. Conclusion
We have surveyed the literature on adversarial contention resolution by delineating results based on key model assumptions, metrics, and algorithm-design choices. The problem of contention resolution remains an active topic of research, and here we briefly discuss some potential directions for future research.
Machine learning (ML) may have an important role to play in the design of contention-resolution algorithms; certainly, the use of ML is being explored in the development of future wireless standards, such as 6G (kato:ten, ). As discussed in Section 3.3.2, Gilbert et al. (gilbert2021contention, ) demonstrate the impact of system-size predictions, which might be offered by a black-box ML module. A natural question is whether there are other useful predictions that could give improved performance, or at least performance that is parameterized by the accuracy of the predictions. This has been done recently for denial-of-service attacks in the traditional client-server model (chakraborty:bankrupting, ). Can predictions about whether a slot will be noisy (or jammed) yield improved robustness?
In a similar vein, incorporating capabilities enabled by new (but existing) wireless technology can yield novel results (e.g., coded networks (bender:coded, )) and deserves further exploration. For instance, the 5G New Radio (NR) ability to dynamically change the width of the frequency band being used can help mitigate congestion or interference on the channel, and allow for power savings (e.g., (abinader:impact, )). Defining a clean abstract model for these capabilities and studying their effects on fundamental problems like contention resolution could be a valuable research focus.
In many wireless systems, collisions have significant cost. For example, in WiFi networks, the entire packet must be sent before the sender learns of success or failure (due to a collision). Under coded networks, multiple packets may be sent simultaneously; therefore, a collision ruins what could otherwise be multiple successful transmissions in a slot. In mobile settings, failure to transmit a packet may mean incurring a significant delay before the sender is next within range of the intended receiver. Therefore, modeling collision cost is another angle that presents some interesting open problems in the area of contention resolution.
Mobility adds a layer of complexity to wireless communication because participants may be constantly changing position, affecting signal strength and interference. This is an aspect that has not received attention in the literature, and integrating mobility into contention resolution might open the door to interesting new problems. Additional costs due to distance traveled or energy expended to overcome attenuation are aspects that may be worthwhile considering.
Adopting different channel models can yield distinct upper and lower bounds compared to the classical model, as we have seen in the case of the SINR model (e.g., (fineman:contention, ; jurdzinski2015cost, )). The power sensitive model explored by De Marco and Kowalski (DBLP:conf/icdcs/MarcoK10, ) and by Tsybakov tsybakov:multiplicity) is another example of a interesting channel model. In the latter case, exploring via experiment the degree to which this model holds in practice may be interesting, as it seems to impart a non-trivial amount of power (e.g., we can count the number of users in $O(1)$ slots rather than ${\texttt{poly}({\log n})}$ slots). Even coarser-grained feedback —rather an exact value—for the number of senders could be useful. For example, in the classical channel, it is hard to tell if noise comes from many users competing for access or from jamming; however, (even rough) estimates could help clarify this.
@@
Energy efficiency is likely to remain an important consideration in wireless systems. The exact limits of trading off between robustness to worst-case noise, energy efficiency, and latency still remain unknown. Can we obtain $o(\log n)$ number of channel-access attempts per packet despite jamming, while retaining asymptotically optimal latency? Related to this aspect, in some wireless standards, transmission power can be increased (and decreased), allowing devices to compensate for signal attenuation. This capability might be useful in overcoming jamming attacks, where good participants incur higher energy costs to transmit at higher power in order to overcome malicious noise. Is it possible to design algorithms in the spirit of (ChangJP19, ; BenderFGY19, ), where the sending costs are grow very slowly with the jamming costs? Alternatively, can the latency be asymptotically improved over existing jamming-resistant results?